Show that;sin(90°-theta)/Cosec(90°-theta)+cos(90°-theta)/sec(90°-theta)=1
1
Expert's answer
2019-11-12T11:18:25-0500

"sin(90^\circ-\theta)=cos(\theta)"

"cos(90^\circ-\theta)=sin(\theta)"

"sec (90^\circ-\theta)=\frac{1}{cos(90^\circ-\theta)}=\frac{1}{sin(\theta)}"

"Cosec (90^\circ-\theta)=\frac{1}{sin(90^\circ-\theta)}=\frac{1}{cos(\theta)}"

"\frac{sin(90^\circ-\theta)}{cosec(90^\circ-\theta)}+\frac{cos(90^\circ-\theta)}{sec(90^\circ-\theta)}=\frac{cos(\theta)}{sec(\theta)}+\frac{sin(\theta)}{cosec(\theta)}=\frac{cos(\theta)}{\frac{1}{cos(\theta)}}+\frac{sin(\theta)}{\frac{1}{sin(\theta)}}="

"=cos^2(\theta)+sin^2(\theta)=1"